OpenStudy (anonymous):

Which measure of central tendency is most appropriate for the data set? Why? 133332333

OpenStudy (unklerhaukus):

\(\color{black}{\{1,3,3,3,3,2,3,3,3\}}\) \(\{1,\color{blue}{3},\color{blue}{3},\color{blue}{3},\color{blue}{3},2,\color{blue}{3},\color{blue}{3},\color{blue}{3}\}\) \({\color{blue} {\text{mode}} \text{ is } }3\) \(\{1,2,3,3,\color{green}3,3,3,3,3\}\) \(\color{green}{\text{median}}\text{ is }3\) \[\color{red}{\frac{\sum\limits_{i=1}^{n}x_i}{n}}=\frac{1+3+3+3+3+2+3+3+3}{9}\]\(\color{red}{\text{mean}}\text{ is } 2\frac 23\)

OpenStudy (anonymous):

MODE is the best

OpenStudy (unklerhaukus):

why is mode the best?

OpenStudy (anonymous):

Mode is the frequent entry that is repeating again and again and here 3 is doing the same..

OpenStudy (unklerhaukus):

how could the mode be better than the median?

OpenStudy (anonymous):

I think I am talking about this particular case..

OpenStudy (unklerhaukus):

in this particular case taking the mode of the data set will yield the same result as taking the median \(\color{black}3\), but the mean has more information, and i think more accurately describes the set of data

OpenStudy (anonymous):

thx :)